# Open-MAGVIT2: An Open-Source Project Toward Democratizing Auto-regressive Visual Generation

### v2

@@ Abstract
We present Open-MAGVIT2, a family of auto-regressive image generation models ranging from 300M to 1.5B. The Open-MAGVIT2 project produces an open-source replication of Google’s MAGVIT-v2 tokenizer, a tokenizer with a super-large codebook (i.e., $2^{18}$ codes), and achieves the state-of-the-art reconstruction performance (1.17 rFID) on ImageNet $256\times 256$. Furthermore, we explore its application in plain auto-regressive models and validate scalability properties. To assist auto-regressive models in predicting with a super-large vocabulary, we factorize it into two sub-vocabulary of different sizes by asymmetric token factorization, and further introduce “next sub-token prediction” to enhance sub-token interaction for better generation quality. We release all models and codes to foster innovation and creativity in the field of auto-regressive visual generation.
@@ 1 Introduction
Large Language Models (LLMs), built upon auto-regressive transformer (Vaswani et al., 2017; OpenAI, 2023; Chowdhery et al., 2022; Touvron et al., 2023), have demonstrated dominance in natural language generation due to the incredible context modeling and scalability. Inspired by this, emergent works introduce auto-regressive models into visual generation (Van Den Oord et al., 2017; Esser et al., 2021; Yu et al., 2022; Lee et al., 2022; Sun et al., 2024). These approaches first utilize a vector quantizer for image tokenization and de-tokenization, then employ an auto-regressive transformer for discrete image token sequence modeling.
@@
In this work, we push forward the auto-regressive visual generation in two folds: 1) Replication of the visual tokenizer: We re-implement the advanced Lookup-Free Quantizer proposed by MAGVIT-v2. To our best knowledge, our open-source replication achieves the closest reconstruction performance stated in MAGVIT-v2 (1.18 vs. 1.15 rFID on ImageNet 128$\times$128) and outperforms all other methods on the hallmark Imagenet benchmark (Deng et al., 2009). Furthermore, we provide the visual tokenizer for text-conditional image generation with improved representational capacity by pretraining it on large-scale general-domain datasets. 2) Integrating a super-large codebook with AR visual generation: Instead of simply following MAGVIT-v2 that leverages the vision-oriented design (i.e., mask generative methods (Chang et al., 2022) for visual synthesis), we seek to exploit the potential of such a large codebook in vanilla auto-regressive generation. To assist auto-regressive models in predicting with a super-large vocabulary, we factorize it into two sub-vocabulary of different sizes by asymmetric token factorization, and further introduce “next sub-token prediction” to enhance sub-token interaction for better generation quality. Our experiments on the standard visual generation dataset ImageNet suggest that, with the powerful tokenizer, the plain auto-regressive model exhibits superiority and scalability.
Table 1: Model configurations of Open-MAGVIT2. We partially follow the scaling rule proposed in the previous works (Sun et al., 2024; Tian et al., 2024).
Model Parameters Inter-Blocks $N$ Intra-Blocks $L$ Widths $w$ Heads $h$
Open-MAGVIT2-B 343M 24 2 1024 16
Open-MAGVIT2-L 804M 36 3 1280 20
Open-MAGVIT2-XL 1.5B 48 4 1536 24
@@ 2.2 Visual Tokenizer
Visual tokenization is fundamentally deemed as the crucial component in multimodal large language models (MLLMs) to understand the visual signal input. The CNN-based encoder-quantizer-decoder architecture first proposed in VQVAE (Van Den Oord et al., 2017) is well adopted as the visual tokenizer, which maps input pixels into discrete representations and reconstructs images from quantized features. Specifically, given an image $\mathcal{I}\in\mathbb{R}^{3\times H\times W}$, the encoder projects it into the feature map $\mathcal{Z}\in\mathbb{R}^{D\times H^{\prime}\times W^{\prime}}$, where $H^{\prime}=H/p,W^{\prime}=W/p,$ and $p$ is the downsample ratio. The quantizer containing a learnable codebook $\mathcal{E}\in\mathbb{R}^{2^{K}\times D}$ then selects the closest entry $\hat{z}\in\mathbb{R}^{D}$ from the codebook for each feature vector $z\in\mathbb{R}^{D}$. And we can use discrete token indices $\mathcal{X}=\{x_{i}\}_{i=1}^{H^{\prime}\times W^{\prime}}$ to represent the continuous feature map $\mathcal{Z}$. For decoding, each code index will be mapped back to the quantized feature vector and input into the decoder for pixel-level image reconstruction.
@@ 3.2 Implementations Details
Auto-regressive Transformer Setup.
3.3 Main Results
Table 2: Model designs and reconstruction performance comparison with the original MAGVIT-v2 on $128\times 128$ ImageNet 50k validation set, following the MAGVIT-v2 paper.
Method Tokens Train LFQ Large Up/Down Deeper Adaptive rFID
@@ 3.3 Main Results
Open-MAGVIT2 $16\times 16$ $128\times 128$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 1.18
MAGVIT2 (Yu et al., 2024a) $16\times 16$ $128\times 128$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 1.15
Table 3: Reconstruction performance of different tokenizers on $256\times 256$ ImageNet 50k validation set. Open-MAGVIT2 achieves SOTA results on different downsampling rates. $\dagger$ specifies that the training is on OpenImages. * denotes that the results are from the direct inference using the model trained with $128\times 128$ resolution without fine-tuning.
@@
Open-MAGVIT2 2D $16\times 16$ 16 $256\times 256$ 262144 1.17 22.64 100%
@@
OmiTokenizer-VQ (Wang et al., 2024) 2D $32\times 32$ 8 $256\times 256$ 8192 1.11 -- --
@@
Open-MAGVIT2 2D $32\times 32$ 8 $128\times 128$ 262144 0.34 27.02 100%
@@
Visual Reconstruction.
Furthermore, to facilitate the development of autoregressive text-conditional image generation, we provide an improved version of our tokenizers by pretraining on large-scale image-text datasets. As illustrated in Fig. 3, our tokenizers achieve state-of-the-art performance compared to concurrent methods such as Cosmos (Agarwal et al., 2025) in terms of zero-shot reconstruction on both ImageNet and COCO datasets. It is worth noting that some recent efforts in residual tokenization (Qu et al., 2024; Han et al., 2024) can achieve better results, but are not listed here because residual techniques are orthogonal and compatible with quantization methods such as VQ, LFQ, and FSQ.
Visual Generation.
@@
Open-MAGVIT2-B 343M 3.08 258.26 0.85 0.51
Open-MAGVIT2-L 804M 2.51 271.70 0.84 0.54
Open-MAGVIT2-XL 1.5B 2.33 271.77 0.84 0.54
Table 5: Zero-shot reconstruction performance on ImageNet 50k validation set and MS-COCO val2017. The tokenizers are trained with large-scale general-domain datasets and aim to serve text-conditional image generation. The results are reported under the same setup. $\dagger$ indicates that LlamaGen loads the model initially trained on Imagenet while the others are training from scratch, i.e., MS-COCO and Imagenet-1k are excluded from training data. Text in gray signifies the metrics adopted from Cosmos (Agarwal et al., 2025) report.
@@
Open-MAGVIT2 LFQ 100M 16 16384 7.93 22.21 0.62 2.55 22.21 0.62
Open-MAGVIT2 LFQ 100M 16 262144 6.76 22.31 0.65 1.67 22.70 0.64
@@
Open-MAGVIT2 LFQ 100M 16 16384 6.65 21.61 0.57 1.39 21.74 0.56
Open-MAGVIT2 LFQ 100M 16 262144 5.10 22.18 0.60 0.78 22.24 0.59
@@ 3.4 Qualitative Results
We present the qualitative results on Imagenet Benchmark in terms of visual reconstruction (see in Fig. 5) and visual generation (see in Fig. 6), respectively.

### v3

@@ Abstract
The Open-MAGVIT2 project produces an open-source replication of Google’s MAGVIT-v2 tokenizer, a tokenizer with a super-large codebook (i.e., $2^{18}$ codes), and achieves the state-of-the-art reconstruction performance on ImageNet and UCF benchmarks. We also provide a tokenizer pre-trained on large-scale data, significantly outperforming Cosmos on zero-shot benchmarks (1.93 vs. 0.78 rFID on ImageNet original resolution). Furthermore, we explore its application in plain auto-regressive models to validate scalability properties, producing a family of auto-regressive image generation models ranging from 300M to 1.5B. To assist auto-regressive models in predicting with a super-large vocabulary, we factorize it into two sub-vocabulary of different sizes by asymmetric token factorization, and further introduce “next sub-token prediction” to enhance sub-token interaction for better generation quality. We release all models and codes to foster innovation and creativity in the field of auto-regressive visual generation.
@@
Table 1: Configurations of visual tokenizers in Open-MAGVIT2 series. We provide both image and video tokenizers, where the former has a pretrained version to facilitate text-conditional image generation.
Tokenizer Training Codebook Size Resolution Temporal Spatial
Data Ratio $p_{t}$ Ratio $p_{s}$
Open-MAGVIT2-I-In1k ImageNet1k 262144 $256\times 256$- $16\times 16$
$128\times 128$- $8\times 8$
Open-MAGVIT2-I-PT Image-text data 16384 $256\times 256$- $16\times 16$
262144 $256\times 256$- $16\times 16$
Open-MAGVIT2-V-UCF UCF-101 262144 $128\times 128$ 4 $8\times 8$
Table 2: Configurations of auto-regressive image generation models in Open-MAGVIT2 series. We partially follow the scaling rule proposed in the previous works (Sun et al., 2024; Tian et al., 2024).
Model Parameters Inter-Blocks $N$ Intra-Blocks $L$ Widths $w$ Heads $h$
Open-MAGVIT2-AR-B 343M 24 2 1024 16
Open-MAGVIT2-AR-L 804M 36 3 1280 20
Open-MAGVIT2-AR-XL 1.5B 48 4 1536 24
@@ 1 Introduction
Large Language Models (LLMs), built upon auto-regressive transformer (Vaswani et al., 2017; OpenAI, 2023; Chowdhery et al., 2022; Touvron et al., 2023), have demonstrated dominance in natural language generation due to the incredible context modeling and scalability. Inspired by this, emergent works introduce auto-regressive models into visual generation (Van Den Oord et al., 2017; Esser et al., 2021; Yu et al., 2022; Lee et al., 2022; Sun et al., 2024). These approaches first utilize a vector quantizer for visual tokenization and de-tokenization, then employ an auto-regressive transformer for discrete visual token sequence modeling.
@@
In this work, we push forward the auto-regressive visual generation in two folds:
  • Replication of the visual tokenizer: We re-implement the advanced Lookup-Free Quantizer proposed by MAGVIT-v2. To our best knowledge, our open-source replication achieves the closest reconstruction performance stated in MAGVIT-v2 (1.18 vs. 1.15 rFID on ImageNet 128$\times$128) and outperforms all other methods on the hallmark Imagenet and UCF benchmarks. Furthermore, we provide the visual tokenizer for text-conditional image generation with improved representational capacity by pretraining it on large-scale general-domain datasets. Our pretrained visual tokenizer achieves state-of-the-art zero-shot reconstruction performance on COCO and ImageNet, surpassing Cosmos (Agarwal et al., 2025) and Llamagen (Sun et al., 2024) significantly (Table 6).
  • Integrating a super-large codebook with AR visual generation: Instead of simply following MAGVIT-v2 that leverages the vision-oriented design (i.e., mask generative methods (Chang et al., 2022) for visual synthesis), we seek to exploit the potential of such a large codebook in vanilla auto-regressive generation. To assist auto-regressive models in predicting with a super-large vocabulary, we factorize it into two sub-vocabulary of different sizes by asymmetric token factorization, and further introduce “next sub-token prediction” to enhance sub-token interaction for better generation quality. Our experiments on the standard visual generation dataset ImageNet suggest that, with the powerful tokenizer, the plain auto-regressive model exhibits superiority and scalability.
Update Log
  • V0.1 (17/06/2024): Release an initial version of Open-MAGVIT2 image tokenizer.
  • V1.0 (09/09/2024): Release a series of Open-MAGVIT2 image tokenizers trained on ImageNet (dubbed as Open-MAGVIT2-I-In1k in the paper) and a family of auto-regressive models ranging from 300M to 1.5B (dubbed as Open-MAGVIT2-AR-B/L/XL in the paper).
  • V1.1 (21/01/2025): Release a pretrained version of Open-MAGVIT2 image tokenizers (dubbed as Open-MAGVIT2-I-PT in the paper).
  • V1.2 (09/02/2025): Release an Open-MAGVIT2 video tokenizer with 4$\times$ temporal downsampling rate (dubbed as Open-MAGVIT2-V-UCF in the paper).
@@ 2.2 Visual Tokenizer
Visual tokenization is fundamentally deemed as the crucial component in multimodal large language models (MLLMs) to understand the visual signal input. The CNN-based encoder-quantizer-decoder architecture first proposed in VQVAE (Van Den Oord et al., 2017) is well adopted as the visual tokenizer, which maps input pixels into discrete representations and reconstructs images from quantized features. Specifically, given an video $\mathcal{V}\in\mathbb{R}^{T\times 3\times H\times W}$ (When $T=1$, the input is an image), the encoder projects it into the feature map $\mathcal{Z}\in\mathbb{R}^{T^{\prime}\times D\times H^{\prime}\times W^{\prime}}$, where $T^{\prime}=T/p_{t},H^{\prime}=H/p_{s},W^{\prime}=W/p_{s},$ and $p_{t}$, $p_{s}$ are the temporal and spatial downsample ratio respectively. The quantizer containing a learnable codebook $\mathcal{E}\in\mathbb{R}^{2^{K}\times D}$ then selects the closest entry $\hat{z}\in\mathbb{R}^{D}$ from the codebook for each feature vector $z\in\mathbb{R}^{D}$. And we can use discrete token indices $\mathcal{X}=\{x_{i}\}_{i=1}^{T^{\prime}\times H^{\prime}\times W^{\prime}}$ to represent the continuous feature map $\mathcal{Z}$. For decoding, each code index will be mapped back to the quantized feature vector and input into the decoder for pixel-level image reconstruction.
@@ 3.1 Dataset and Metrics
Video Generation.
Following (Yu et al., 2024a), we train our video tokenizer on UCF-101 (Soomro, 2012). We use 17-frame video clips with a spatial resolution of 128 $\times$ 128 for both training and evaluation. The reconstruction-FVD, denoted as rFVD (Unterthiner et al., 2018) is adopted as the main metrics for quantifying the soundness of reconstructed videos.
@@ 3.2 Implementations Details
Table 3: Model designs and reconstruction performance comparison with the original MAGVIT-v2 on $128\times 128$ ImageNet 50k validation set (rFID) and UCF-101 (rFVD).
Method Tokens Train LFQ Large Up/Down Deeper Adaptive rFID rFVD
@@
Open-MAGVIT2 $16\times 16$ $128\times 128$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 1.18 16.0
MAGVIT2 (Yu et al., 2024a) $16\times 16$ $128\times 128$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 1.15 8.6
@@
Open-MAGVIT2-I-In1k 2D $16\times 16$ 16 $256\times 256$ 262144 1.17 22.64 100%
@@
OmiTokenizer-VQ (Wang et al., 2024b) 2D $32\times 32$ 8 $256\times 256$ 8192 1.11 -- --
@@
Open-MAGVIT2-I-In1k 2D $32\times 32$ 8 $128\times 128$ 262144 0.34 27.02 100%
@@
Video Tokenizer
TATS (Ge et al., 2022) 2D $4\times 16\times 16$ 8 $128\times 128$ 16384 162 -- --
MAGVIT (Yu et al., 2023) 2D $4\times 16\times 16$ 8 $128\times 128$ 1024 25 -- --
SweetTokenizer (Tan et al., 2024) 1D 256 + 1024 -- $256\times 256$ 10481 + 11139 44 -- --
LARP-L (Wang et al., 2024a) 1D 1024 -- $128\times 128$ 8192 24 -- --
LARP-L-Long (Wang et al., 2024a) 1D 1024 -- $128\times 128$ 8192 20 -- --
SweetTokenizer†† (Tan et al., 2024) 1D 5120 -- $256\times 256$ 10481 + 11139 18 -- --
Open-MAGVIT2-V-UCF 2D $5\times 16\times 16$ 8 $128\times 128$ 262144 16 -- 100
@@
Open-MAGVIT2-AR-B 343M 3.08 258.26 0.85 0.51
Open-MAGVIT2-AR-L 804M 2.51 271.70 0.84 0.54
Open-MAGVIT2-AR-XL 1.5B 2.33 271.77 0.84 0.54
Table 6: Zero-shot reconstruction performance on ImageNet 50k validation set and MS-COCO val2017. The tokenizers are trained with large-scale general-domain datasets and aim to serve text-conditional image generation. The results are reported under the same setup for fair comparison (text in gray signifies the results directly from Cosmos (Agarwal et al., 2025) report). $\dagger$ indicates that LlamaGen loads the model initially trained on Imagenet while the others are training from scratch, i.e., MS-COCO and Imagenet-1k are excluded from training data.
@@
Open-MAGVIT2-I-PT LFQ 100M 16 16384 7.93 22.21 0.62 2.55 22.21 0.62
Open-MAGVIT2-I-PT LFQ 100M 16 262144 6.76 22.31 0.65 1.67 22.70 0.64
@@
Open-MAGVIT2-I-PT LFQ 100M 16 16384 6.65 21.61 0.57 1.39 21.74 0.56
Open-MAGVIT2-I-PT LFQ 100M 16 262144 5.10 22.18 0.60 0.78 22.24 0.59
Video Tokenizer Setup.
We extend the image version of tokenizer following (Yu et al., 2024a) where the temporally casual 3D convolution serves as the main design. Simultaneously, we inflate the image tokenizer trained at 128 $\times$ 128 imagenet for video modeling. Similarly, we adopt 3D PatchGAN as the discriminator. The model is trained with the following settings: an initial $41e-$ 4 learning rate with a cosine annealing scheduler, an Adam Optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.9$, a total $128$ batch size with 2500 epochs, the losses are the same as the image tokenizer and $64$$\times$ GPU / NPU with Pytorch.
Auto-regressive Transformer Setup.
3.3 Main Results
Visual Reconstruction.
To facilitate the development of autoregressive text-conditional image generation, we provide an improved version of our tokenizers by pretraining on large-scale image-text datasets. As illustrated in Fig. 4, our tokenizers achieve state-of-the-art performance compared to concurrent methods such as Cosmos (Agarwal et al., 2025) in terms of zero-shot reconstruction on both ImageNet and COCO datasets. It is worth noting that some recent efforts in residual tokenization (Qu et al., 2024; Han et al., 2024) can achieve better results, but are not listed here because residual techniques are orthogonal and compatible with quantization methods such as VQ, LFQ, and FSQ.
Furthermore, we extend our image tokenizer to the video version. As depicted in Tab. 3, The reconstruction result gap on UCF-101 between MAGVIT-v2 and us is narrow. It is worth noting that Open-MAGVIT2 achieves a competitive performance among previous video tokenizers (see in Tab. 4)
Visual Generation.
@@ 3.4 Qualitative Results
We present the qualitative results on Imagenet Benchmark in terms of visual reconstruction (see in Fig. 6) and visual generation (see in Fig. 7), respectively. The video reconstruction visualization (see in Fig. 5) are on UCF-101.
@@ 4 Related Works
Figure 5: Visualization of Open-MAGVIT2 video reconstruction. (a) indicates the original videos while (b) specifies the reconstruction videos.